# Simulation of chiral motion of excitation within the ground-state manifolds of neutral atoms

### v1

@@ IV.3 Coherence time of the system
$H_{I}^{(2)}$ $=$ $\Omega_{2}^{\beta_{2}}e^{-i\delta_{2}t}|r_{1}\rangle\langle g_{1}|+\Omega_{2}^{\alpha_{2}}e^{-i\delta_{2}t}|r_{2}\rangle\langle g_{2}|$ (14)
$+\Omega_{p}\sum_{j=2}^{3}|r_{j}\rangle\langle e_{j}|e^{-i\Delta t}+|r_{1}r_{2}\rangle\langle r_{1}r_{2}|+{\rm H.c.},$
@@
Hao-Yuan Tang: Data curation (equal); Writing-original draft. Xiao-Xuan Li: Data curation (equal). Jia-Bin You: Writing-review & editing (equal); Supervision (equal); Funding acquisition (equal). Xiao-Qiang Shao: Conceptualization (lead); Writing-review & editing (equal); Supervision (equal); Funding acquisition (equal).
@@ Appendix A The details of Hamiltonian (3)
$H_{I}$ $=$ $(\Omega_{1}^{\alpha_{1}}e^{-i\delta_{1}t}+\Omega_{2}^{\beta_{2}}e^{-i\delta_{2}t})|r_{1}\rangle\langle g_{1}|+(\Omega_{2}^{\alpha_{2}}e^{-i\delta_{2}t}$ (15)
@@
$+\Omega_{p}\sum_{j=1}^{3}|r_{j}\rangle\langle e_{j}|e^{-i\Delta t}+\frac{1}{2}\sum_{i\neq j}\mathcal{U}_{i,j}|r_{i}r_{j}\rangle\langle r_{i}r_{j}|+{\rm H.c.}.$
Considering a rotating frame with respect to $U=\exp[-i\sum_{i\neq j}\mathcal{U}_{i,j}/2|r_{i}r_{j}\rangle\langle r_{i}r_{j}|t]$, and $\mathcal{U}_{i,j}=\Delta$. The Hamiltonian satisfies $H_{I}^{{}^{\prime}}=i\dot{U}^{{\dagger}}U+U^{{\dagger}}H_{I}U$, then the Eq. (15) can be reformulated as:
@@
To verify the accuracy of our effective model, we compared the dynamics of the system governed by the full Hamiltonian (15) with those described by Eq. (18), as shown in Fig. 7. We found that these two cases match very well.
Figure 7: The chiral motion of atomic excitation among the system governed by the full Hamiltonian (15) and Eq. (18), respectively. The parameters are $(\Omega_{1},\Omega_{2},\Omega_{3},\Omega_{p})/(2\pi)=(0.2,0.2,0.4,4)$ MHz, $(\alpha_{1},\alpha_{j\neq 1},\beta_{j})=(\pi/{2},0,0)$, and $(\delta_{1},\delta_{2},\delta_{3},\Delta)/(2\pi)=(-4,4,8,200)$ MHz.

### v2

@@ IV.3 Coherence time of the system
$H_{I}^{(2)}$ $=$ $\Omega_{2}^{\beta_{2}}e^{-i\delta_{2}t}|r_{1}\rangle\langle g_{1}|+\Omega_{2}^{\alpha_{2}}e^{-i\delta_{2}t}|r_{2}\rangle\langle g_{2}|$
$+\Omega_{p}\sum_{j=2}^{3}|r_{j}\rangle\langle e_{j}|e^{-i\Delta t}+{\rm H.c.}+\mathcal{U}_{1,2}|r_{1}r_{2}\rangle\langle r_{1}r_{2}|,$
@@
Hao-Yuan Tang: Data curation (equal); Writing-original draft (lead). Xiao-Xuan Li: Data curation (equal). Jia-Bin You: Writing-review & editing (equal); Supervision (equal); Funding acquisition (equal). Xiao-Qiang Shao: Conceptualization (lead); Writing-review & editing (equal); Supervision (equal); Funding acquisition (equal).
@@ Appendix A The details of Hamiltonian (3)
$H_{I}$ $=$ $(\Omega_{1}^{\alpha_{1}}e^{-i\delta_{1}t}+\Omega_{2}^{\beta_{2}}e^{-i\delta_{2}t})|r_{1}\rangle\langle g_{1}|+(\Omega_{2}^{\alpha_{2}}e^{-i\delta_{2}t}$
@@
$+\Omega_{p}\sum_{j=1}^{3}|r_{j}\rangle\langle e_{j}|e^{-i\Delta t}+{\rm H.c.}+\frac{1}{2}\sum_{i\neq j}\mathcal{U}_{i,j}|r_{i}r_{j}\rangle\langle r_{i}r_{j}|.$
Considering a rotating frame with respect to $U=\exp[-i\sum_{i\neq j}\mathcal{U}_{i,j}/2|r_{i}r_{j}\rangle\langle r_{i}r_{j}|t]$, and $\mathcal{U}_{i,j}=\Delta$. The Hamiltonian satisfies $H_{I}^{{}^{\prime}}=i\dot{U}^{{\dagger}}U+U^{{\dagger}}H_{I}U$, then the Eq. (A) can be reformulated as:
@@
To verify the accuracy of our effective model, we compared the dynamics of the system governed by the full Hamiltonian (A) with those described by Eq. (18), as shown in Fig. 7. We found that these two cases match very well.
Figure 7: The chiral motion of atomic excitation among the system governed by the full Hamiltonian (A) and Eq. (18), respectively. The parameters are $(\Omega_{1},\Omega_{2},\Omega_{3},\Omega_{p})/(2\pi)=(0.2,0.2,0.4,4)$ MHz, $(\alpha_{1},\alpha_{j\neq 1},\beta_{j})=(\pi/{2},0,0)$, and $(\delta_{1},\delta_{2},\delta_{3},\Delta)/(2\pi)=(-4,4,8,200)$ MHz.